# Multiply-Robust Causal Change Attribution

### v1

@@ 2.4 Estimation
In this section, we describe how to implement the multiply-robust estimating equation (MR) in practice.
@@
The nested regression functions $\gamma_{k}(\bar{\bm{X}}_{k})$ can be estimated by any parametric or non-parametric ML regression algorithm, including LASSO, ridge, random forests, neural networks, boosting, or any ensemble method combining multiple of these. We work recursively, starting with an estimator $\hat{\gamma}_{K}(\bar{\bm{X}}_{K})$ for the regression of $h(Y)$ on $\bar{\bm{X}}_{K}:=\bm{X}$, using the the data from sample $c_{K+1}\in\{0,1\}$ (i.e., the distribution we want to fix for $P_{Y\mid\bm{X}}$). Next, we obtain $\hat{\gamma}_{K-1}(\bar{\bm{X}}_{K-1})$ by regressing $\hat{\gamma}_{K}(\bar{\bm{X}}_{K})$ on $\bar{\bm{X}}_{K-1}$, using the the data from sample $c_{K}\in\{0,1\}$ (i.e., the distribution we want to fix for $P_{X_{K}\mid\mathrm{PA}_{K}}$). We proceed this way until up to $\hat{\gamma}_{1}(X_{1})$.
@@
Finally, we build an estimator of the counterfactual parameter $\theta^{\bm{c}}$ by replacing $g(\bar{\bm{X}}_{k})$ with $\hat{\gamma}_{k}(\bar{\bm{X}}_{k})$ and $a(\bar{\bm{X}}_{k})$ with $\hat{\alpha}_{k}(\bar{\bm{X}}_{k})$ in a sample analog of equation (MR):
$\hat{\theta}^{\bm{c}}:=\widehat{\mathrm{E}}_{(c_{1})}[\hat{\gamma}_{1}(X_{1})]+\sum_{k=1}^{K}\widehat{\mathrm{E}}_{(c_{k+1})}[\hat{\alpha}_{k}(\bar{\bm{X}}_{k})\{\hat{\gamma}_{k+1}(\bar{\bm{X}}_{k+1})-\hat{\gamma}_{k}(\bar{\bm{X}}_{k})\}],$ (EST)
where we use $\widehat{\mathrm{E}}_{(t)}[f(W)]=n_{t}^{-1}\sum_{i:T_{i}=t}f(W_{i})$ to denote the sample average on sample $t\in\{0,1\}$.
@@
To avoid overfitting bias, our recommendation is that the data used to learn the unknown functions $\gamma_{k}(\bar{\bm{X}}_{k})$, $\alpha_{k}(\bar{\bm{X}}_{k})$ are not re-used to estimate the main parameter $\theta^{\bm{c}}$. When the dataset is large, different random subsamples can be used. In medium-sized datasets, a good alternative is to use cross-fitting (see, for example, Newey & Robins, 2018 for a definition and theoretical advantatges of cross-fitting). ∎
@@ 2.5 Large-Sample Inference
In this section, we show consistency and asymptotic normality of our estimator $\hat{\theta}^{\bm{c}}$, and explain how to perform large-sample inference on the true parameter $\theta^{\bm{c}}$. First, we notice that the estimator in (EST) is the sum of two averages over different samples:
@@ 2.6 Quantifying the Contribution of Each Causal Mechanism
As discussed in the literature (Budhathoki et al., 2021), the effect of changing the one causal mechanism may depend on which other mechanisms have already been changed; Shapley values take that into account, whereas the method along a causal path does not. Whether this is a relevant concern depends on the application: in some cases, the impact of a given explanatory variable may be approximately the same regardless of the distribution of other explanatory variables. On the other hand, causal attribution along a causal path requires estimating $\theta^{\bm{c}}$ for only $K$ change vectors $\bm{c}$ (and, by the simplification in C.1, each requires training only one regressor and one classifier), whereas computing Shapley values requires estimating $\theta^{\bm{c}}$ for $2^{K+1}$ change vectors $\bm{c}$ (each of which could require training up to $2K$ machine learners). Some computationally efficient approximations to $\mathrm{SHAP}_{k}$ exist (e.g, Kolpaczki et al., 2023).
Let $\widehat{\mathrm{SHAP}}_{k}$ and $\widehat{\mathrm{PATH}}_{k}$ be estimators of $\mathrm{SHAP}_{k}$ and $\mathrm{PATH}_{k}$, respectively, that replace the true $\theta^{\bm{c}}$ with the multiply-robust estimator $\hat{\theta}^{\bm{c}}$ in (EST). The following corollary gives a useful result for large-sample inference on the causal attribution parameters:
@@
This is a consequence of $\widehat{\mathrm{SHAP}}_{k}$ and $\widehat{\mathrm{PATH}}_{k}$ being finite linear combinations of $\hat{\theta}^{\bm{c}}$. The asymptotic variance could be computed explicitly, although this might be cumbersome, especially for Shapley values, since it requires the $2^{K+1}\times 2^{K+1}$ covariance matrix of $\hat{\theta}^{\bm{c}}$ for $\bm{c}\in\{0,1\}^{K+1}$. As an alternative, the multiplier bootstrap of Belloni et al. (2017) can be used. We describe the procedure in Appendix D. An advantage of the multiplier approach is that each bootstrap replication does not require re-training the ML estimators of the regression function and weights.
@@ 3.1 Monte-Carlo Simulations
We now present simulation evidence on the performance of our method. Our simulation design is based on the causal model of 2.5.

### v4

@@ 2.4 Estimation
In this section, we describe how to implement the multiply-robust estimating equation (1) in practice.
@@
The nested regression functions $\gamma_{k}(\bar{\bm{X}}_{k})$ can be estimated by any parametric or non-parametric ML regression algorithm, including LASSO, ridge, random forests, neural networks, boosting, or any ensemble method combining multiple of these. We work recursively, starting with an estimator $\hat{\gamma}_{K}(\bar{\bm{X}}_{K})$ for the regression of $h(Y)$ on $\bar{\bm{X}}_{K}:=\bm{X}$, using the the data from sample $c_{K+1}\in\{0,1\}$ (i.e., the distribution we want to fix for $P_{Y\mid\bm{X}}$). Next, we obtain $\hat{\gamma}_{K-1}(\bar{\bm{X}}_{K-1})$ by regressing $\hat{\gamma}_{K}(\bar{\bm{X}}_{K})$ on $\bar{\bm{X}}_{K-1}$, using the the data from sample $c_{K}\in\{0,1\}$ (i.e., the distribution we want to fix for $P_{X_{K}\mid\mathrm{PA}_{K}}$). We proceed this way up until we obtain $\hat{\gamma}_{1}(X_{1})$.
@@
Finally, we build an estimator of the counterfactual parameter $\theta^{\bm{c}}$ by replacing $g(\bar{\bm{X}}_{k})$ with $\hat{\gamma}_{k}(\bar{\bm{X}}_{k})$ and $a(\bar{\bm{X}}_{k})$ with $\hat{\alpha}_{k}(\bar{\bm{X}}_{k})$ in a sample analog of equation (1):
$\hat{\theta}^{\bm{c}}:=\widehat{\mathrm{E}}_{(c_{1})}[\hat{\gamma}_{1}(X_{1})]\\ +\sum_{k=1}^{K}\widehat{\mathrm{E}}_{(c_{k+1})}[\hat{\alpha}_{k}(\bar{\bm{X}}_{k})\hat{e}_{k}(\bar{\bm{X}}_{k+1})],$ (2)
where $\hat{e}_{k}(\bar{\bm{X}}_{i,k+1}):=\hat{\gamma}_{k+1}(\bar{\bm{X}}_{k+1})-\hat{\gamma}_{k}(\bar{\bm{X}}_{k})$ for $k=1,\ldots,K$, and we use $\widehat{\mathrm{E}}_{(t)}[f(W)]=n_{t}^{-1}\sum_{i:T_{i}=t}f(W_{i})$ to denote the sample average on sample $t\in\{0,1\}$.
@@
To avoid overfitting bias, our recommendation is that the data used to learn the unknown functions $\gamma_{k}(\bar{\bm{X}}_{k})$, $\alpha_{k}(\bar{\bm{X}}_{k})$ are not re-used to estimate the main parameter $\theta^{\bm{c}}$. When the dataset is large, different random subsamples can be used. In medium-sized datasets, a good alternative is to use cross-fitting (see, for example, Newey & Robins, 2018). ∎
@@ 2.5 Large-Sample Inference
In this section, we show consistency and asymptotic normality of our estimator $\hat{\theta}^{\bm{c}}$, and explain how to perform large-sample inference on the true parameter $\theta^{\bm{c}}$. First, we notice that the estimator in (2) is the sum of two averages over different samples:
@@ 2.6 Quantifying the Contribution of Each Causal Mechanism
As discussed in the literature (Budhathoki et al., 2021), the effect of changing the one causal mechanism may depend on which other mechanisms have already been changed; Shapley values take that into account, whereas the method along a causal path does not. Whether this is a relevant concern depends on the application: in some cases, the impact of a given explanatory variable may be approximately the same regardless of the distribution of other explanatory variables. On the other hand, causal attribution along a causal path requires estimating $\theta^{\bm{c}}$ for only $K$ change vectors $\bm{c}$ (and, by the simplification in D.1, each requires training only one regressor and one classifier), whereas computing Shapley values requires estimating $\theta^{\bm{c}}$ for $2^{K+1}$ change vectors $\bm{c}$ (each of which could require training up to $2K$ machine learners). Some computationally efficient approximations to $\mathrm{SHAP}_{k}$ exist (e.g, Kolpaczki et al., 2023).
Let $\widehat{\mathrm{SHAP}}_{k}$ and $\widehat{\mathrm{PATH}}_{k}$ be estimators of $\mathrm{SHAP}_{k}$ and $\mathrm{PATH}_{k}$, respectively, that replace the true $\theta^{\bm{c}}$ with the multiply-robust estimator $\hat{\theta}^{\bm{c}}$ in (2). The following corollary gives a useful result for large-sample inference on the causal attribution parameters:
@@
This is a consequence of $\widehat{\mathrm{SHAP}}_{k}$ and $\widehat{\mathrm{PATH}}_{k}$ being finite linear combinations of $\hat{\theta}^{\bm{c}}$. The asymptotic variance could be computed explicitly, although this might be cumbersome, especially for Shapley values, since it requires the $2^{K+1}\times 2^{K+1}$ covariance matrix of $\hat{\theta}^{\bm{c}}$ for $\bm{c}\in\{0,1\}^{K+1}$. As an alternative, the multiplier bootstrap of Belloni et al. (2017) can be used. We describe the procedure in Appendix E. An advantage of the multiplier approach is that each bootstrap replication does not require re-training the ML estimators of the regression function and weights.
@@ 3 Empirical Results
In this section, we discuss two sets of Monte-Carlo simulation results and a real-data application of our method to understanding the gender wage gap. The code for the simulations and the empirical application is available at https://github.com/victor5as/mr_causal_attribution.
@@ 3.1 Monte-Carlo Simulations
We now present simulation evidence on the performance of our method. Our first simulation design is based on the causal model of 2.6, with DAG $X_{1}\rightarrow X_{2}$, $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$. We give details about the data-generating process in Appendix F.
Table 1 presents the results of our simulation (over 1,000 draws) for different choices of learners for the regression and the weights. 6(a) shows the Mean Absolute Error (MAE) for a scenario where both the regression and the weights are correctly specified parametric learners: OLS with quadratic terms for the regressions and logistic classifier with quadratic terms to build the weights. [^2] This is the correct specification for the weights, since the log-likelihood ratio between two Gaussian distributions with possibly different variances is a quadratic polynomial. Comparing our multiply-robust (MR) estimator to using regression or re-weighting only, we can see that its performance is statistically indistinguishable from the best of the two methods (regression).
Tables 6(b) and 6(c) show what happens when either the weights or the regressions are misspecified, while the specification of the other component of the model is still correct. As a particular form of misspecification, we omit the quadratic terms from either the OLS regression or the logistic classifier. In either case, we observe that MR performs as well as the correctly-specified method, while the misspecified method exhibits much higher errors. This illustrates the multiple robustness property of our estimator. When both the weights and the regression are misspecified parametric models, as in 6(d), our theoretical guarantees no longer apply. However, we still show these results for completeness. Although it seems to be the case that the error in Shapley values achieved by the MR methods is about as low as the best of the two other methods, it is significantly larger than the correctly-specified benchmark of Figure 6(a).
Next, we turn our attention to flexible, non-parametric learners for the unknown regression functions and weights, to mimic a more realistic situation in which we do not want to make strong parametric assumptions about these. First, we consider neural networks (Figure 6(e)). We use MLPRegressor and MLPClassifier from the Python library sklearn. We use the default settings for all hyperparameters, except that we allow for early stopping and use 3 hidden layers of 100 neurons each. Second, we consider gradient boosting (Figure 6(f)), using GradientBoostingRegressor and GradientBoostingClassifier from sklearn with the default settings. In both cases, we calibrate the predicted probabilities after classification with CalibratedClassifierCV also from sklearn. The results clearly illustrate 2.11: the bias of the MR non-parametric estimator will, in general, vanish faster than the convergence rates of regression-only or re-weighting-only methods. As such, we see that in general MR has equal or lower error than the best of the other two methods. Moreover, the MAE is often statistically indistinguishable from that of the correctly-specified parametric benchmark 6(a).
Design 2
In a second set of experiments, we consider the effect of increasing hte number of explanatory variables $K$. We consider a line DAG, $X_{1}\rightarrow X_{2}\rightarrow\cdots\rightarrow X_{K}\rightarrow Y$. We give more details about the data-generating process in Appendix F. We use LASSO to learn the regression functions, and Logistic Regression with $\ell_{1}$ penalty (Logit-LASSO) for the weights. In both cases, the penalty is selected by cross-validation.
Table 2 shows the average (across 500 simulations) of the worst-case (across parameters) absolute error, that is, the average of $\max_{k=1,\ldots,K+1}|\mathrm{PATH}_{k}-\widehat{\mathrm{PATH}}_{k}|$. The pattern is almost identical if we look at the Mean Absolute Error instead. The MR estimator performs significantly better than the other two methods, even though the regression and the weights are correctly specified (i.e., they are truly linear/logistic and sparse).
3.2 Real-World Application: Gender Wage Gap
EducationOccupationWage
Figure 4: DAG for the gender wage gap application.
We assume the same causal graphical model as Budhathoki et al. (2021), which is captured by the DAG in Figure 4. We randomly split the data into a training set used to estimate the regression function and weights (40%), a calibration set to calibrate the predicted probabilities from our classification algorithm (10%), and an evaluation set used to obtain the main estimates of the $\hat{\theta}^{\bm{c}}$ parameters (50%). To estimate the regression and the weights, we use HistGradientBoostingRegressor and HistGradientBoostingClassifier from sklearn in Python. We calibrate the probabilities by isotonic regression.
We plot our estimated Shapley values are in Figure 5. The contribution of education is estimated to be positive and significant. In contrast, the distribution of occupation given education has a significant negative effect, slightly larger in magnitude than the contribution of education, but of opposite sign. Most of the gender wage gap is therefore explained by differences in the distribution of wages given education and occupation. This is consistent with previous findings in the literature. For example, Chernozhukov et al. (2018b), documents the same fact for the UK.
[…]
4 Conclusions
Finally, we suggest one direction for future research. The causal interpretation of the change attribution parameters relies on an assumption of no unobserved confounding. The sensitivity bounds in Chernozhukov et al. (2022) could be adapted as a way to test the robustness of a causal change attribution study to unobserved confounding.
Acknowledgements
The authors would like to thank Victor Chernozhukov, Dominik Janzing, Eric Tchetgen-Tchetgen, and seminar audiences at Amazon and MIT for providing helpful feedback. We also would like to thank Patrick Blöbaum for help with code integration in DoWhy.
Impact Statement
This paper presents work whose goal is to advance the fields of Causal Inference and Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
The “causal” interpretation of our change attribution measures is grounded on certain assumptions about the Directed Acyclic Graph (DAG) that codifies causal dependence among the variables in the data. We are precise about these assumptions (Section 2.1), and discuss some ways to relax them in future extensions (Section 4). Our theoretical results also rely on regularity conditions, which we state formally in the Appendix and explain in more intuitive terms in the main body of the paper. Practitioners need to be mindful that, if these assumptions do not hold, our proposed algorithm is not guaranteed to lead to accurate results.
Appendix A Proofs
A.1 Proofs of Section 2.2
Proof of Lemma 2.1.
We begin by assuming $P^{(1)}_{X}\ll P^{(0)}_{X}$, that is, $P^{(1)}_{X}$ is absolutely continuous with respect to $P^{(0)}_{X}$. We give a formal definition of absolute continuity in A.1 (Weak Overlap), which we will present in the proof of the general identification result in Theorem 2.5. We also discuss this assumption in 2.4 in the main text. By the Radon-Nykodim Theorem (Billingsley, 1995), the Radon-Nykodim (RN) derivative $\alpha(X):=\mathrm{d}P^{(1)}_{X}/\mathrm{d}P^{(0)}_{X}(X)$ exists, is almost-surely unique. Furthermore, we assume $\mathrm{E}_{(0)}[Y^{2}]<\infty$ and $\mathrm{E}_{(0)}[\alpha(X)^{2}]<\infty$, which is a version of A.2 (Moments). Assumptions A.1 and A.2, combined with the Cauchy-Schwartz inequality, imply that any function $g(X)$ such that $\mathrm{E}_{(0)}[g(X)^{2}]<\infty$ will be $P_{X}^{(1)}$-integrable function $g(X)$, and moreover:
$\mathrm{E}_{(1)}[\gamma(X)]=\int\gamma(x)\mathrm{d}P^{(1)}_{X}(x)=\int\left\{\int y\,\mathrm{d}P^{(0)}_{Y\mid X}(y\mid x)\right\}\mathrm{d}P^{(1)}_{X}(x)=\int y\,\mathrm{d}P^{(0)}_{Y\mid X}(y\mid x)\mathrm{d}P^{(1)}_{X}(x),$
which shows (REG).
which shows (REW). ∎
Proof of Lemma 2.3.
Consider first the case of $g(X)=\gamma(X)$. In that case, by definition of $\gamma(X)$, $\mathrm{E}[e(Y,X)\mid X]=\mathrm{E}[Y-\gamma(X)\mid X]=0$. By the Law of Iterated Expectations, $\mathrm{E}_{(0)}[a(X)e(Y,X)]=0$ for any $a(X)$ such that $\mathrm{E}_{(0)}[a(X)^{2}]<\infty$, and so (DR) and (REG) give:
$\mathrm{E}_{(1)}[\gamma(X)]+\mathrm{E}_{(0)}[a(X)e(Y,X)]=\mathrm{E}_{(1)}[\gamma(X)]=\theta^{\langle 1,0\rangle}.$
A.2 Proofs of Section 2.3
Proof of Theorem 2.5.
We assume the following regularity conditions:
Assumption A.1 (Weak Overlap).
Recall that, for two measures $\mu,\nu$ on a $\sigma$-algebra $\mathcal{B}$, we say that $\nu\ll\mu$ ($\nu$ is absolutely continuous with respect to $\mu$) whenever $\mu(B)=0$ implies $\nu(B)=0$ for $v\in\mathcal{B}$ (Billingsley, 1995). If $\nu\ll\mu$ and $\mu\ll\nu$, we say that $\mu$ and $\nu$ are mutually absolutely continuous. Absolute continuity guarantees that the RN derivatives that go into the weight exist. Intuitively, we are requiring that the distribution of $X_{k}\mid\mathrm{PA}_{k}$ has the same support in samples 0 and 1. This is essential both for regression methods (avoids extrapolating the regression function) and for re-weighting methods (guarantees that we can find observations that will be similar enough to the other sample), as discussed in 2.4.
Assumption A.2 (Moments).
Step 1: Initial Case
as we wanted to show.
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1})]+\mathrm{E}_{(c_{2})}[\alpha_{1}(X_{1})g_{2}(X_{1},X_{2})]-\mathrm{E}_{(c_{2})}[\alpha_{1}(X_{1})g_{1}(X_{1})]=\int g_{2}(x_{1},x_{2})\mathrm{d}P^{(c_{2})}_{X_{2}\mid X_{1}}(x_{2}\mid x_{1})\mathrm{d}P^{(c_{1})}_{X_{1}}(x_{1}),$
as we wanted to show.
Step 2: Induction
Suppose that for some $2\leq j\leq K$, it holds that:
We next show that, if
[…]
(but not necessarily both), we have:
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1})]+\sum_{k=1}^{j-1}\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})e_{k}(\bar{\bm{X}}_{k+1})]+\mathrm{E}_{(c_{j+1})}[\alpha_{j}(\bar{\bm{X}}_{j})g_{j+1}(\bar{\bm{X}}_{j+1})]-\mathrm{E}_{(c_{j+1})}[\alpha_{j}(\bar{\bm{X}}_{j})g_{j}(\bar{\bm{X}}_{j})]\\ =\mathrm{E}_{(c_{j+1})}[\alpha_{j}(\bar{\bm{X}}_{j})\gamma_{j}(\bar{\bm{X}}_{j})].$
Finally, property (RN) and (*A.2) give the desired result.
Step 3: Final Step
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1})]+\sum_{k=1}^{K}\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k})e_{k}(\bar{\bm{X}}_{k+1})]=\int h(y)\mathrm{d}P^{(c_{K+1})}_{Y\mid\bm{X}}(y\mid\bm{x})\prod_{k=1}^{K}\mathrm{d}P^{(c_{k})}_{X_{k}\mid\bar{\bm{X}}_{k-1}}(x_{k}\mid\bar{\bm{X}}_{k-1}).$
A.3 Proofs of Section 2.5
Proof of Theorem 2.10.
Assumption A.3 (Strong Overlap).
Assumption A.4 (Higher-order Moments).
We assume:
  1. (i)
    For some $q>2$ and all $k=1,\ldots,K$, $\mathrm{E}_{(c_{k})}[{\gamma}_{k}(\bar{\bm{X}}_{k})^{q}]\leq C$, $\mathrm{E}_{(c_{k})}[\hat{\gamma}_{k}(\bar{\bm{X}}_{k})^{q}]\leq C$, $\mathrm{E}_{(c_{k+1})}[{\gamma}_{k}(\bar{\bm{X}}_{k})^{q}]\leq C$, and $\mathrm{E}_{(c_{k+1})}[\hat{\gamma}_{k}(\bar{\bm{X}}_{k})^{q}]\leq C$.
  2. (ii)
  3. (iii)
Assumption A.5 (Estimation Rates).
  1. (i)
  2. (ii)
Appendix B Structural Equations and Potential Outcomes
$Y$ $\leftarrow g_{Y}(T,X,\varepsilon_{Y}),$
$X$ $\leftarrow g_{X}(T,\varepsilon_{X}),$
$T$ $\leftarrow\varepsilon_{T},$
$\mathrm{IE}_{0,1}:=\mathrm{E}[Y(0,X(1))]-\mathrm{E}[Y(0)].$
In words, this is the expected effect of holding $T=0$ constant, and changing $X$ to whatever value it would have attained if $T$ had been set to 1, $X(1)$. As long as there are no open backdoor paths relative to $T\rightarrow X$ and $(T,X)\rightarrow Y$, [^3] More generally, we could allow for pre-treatment covariates $W$ that block those backdoor paths. We discuss this generalization in Appendix C.
where $\theta^{\langle 0,0\rangle}=\mathrm{E}_{(0)}[Y]$ can be estimated directly by taking the average of $Y$ in sample 0, and $\theta^{\langle 1,0\rangle}$ can be identified as described in Section 2.2. The mediation formula (Pearl, 2009, equation (4.18)) is capturing the same idea as identification by regression (REG).
Appendix C Identification with Pre-Treatment Covariates
As discussed in Appendix B, under the assumption that $T$ is a randomly assigned treatment, the causal change attribution problem has an interpretation in terms of mediation Pearl (2009, §4.5). Sometimes, random assignment of $T$ may be implausible, but the researcher may have access to a set of pre-treatment covariates $\bm{W}$ such that an unconfoundedness assumption holds conditional on $\bm{W}$. It is still possible to get multiple robustness in that case, with two adjustments: (i) the regression and weights need to control for $\bm{W}$, and (ii) an additional debiasing term is required, to account for the adjustment on covariates. We explain how to do it in this section.
First, notice that the multiply-robust formula given in Theorem 2.5 also holds conditional on covariates, as formalized in the following result.
Proposition C.1 (Identification with Covariates, Conditional Version).
Let $g_{k}(\bar{\bm{X}}_{k},\bm{W})$, $a_{k}(\bar{\bm{X}}_{k},\bm{W})$ be any candidate functions such that $\mathrm{E}_{(c_{k+1})}[g_{k}(\bar{\bm{X}}_{k},\bm{W})^{2}]<\infty$, $\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k},\bm{W})^{2}]<\infty$ for $k=1,\ldots,K$. Consider the following estimating equation:
$\mathrm{E}_{(c_{1})}[g_{1}(X_{1},\bm{W})\mid\bm{W}]+\sum_{k=1}^{K}\mathrm{E}_{(c_{k+1})}[a_{k}(\bar{\bm{X}}_{k},\bm{W})e_{k}(\bar{\bm{X}}_{k+1},\bm{W})\mid\bm{W}],$ (MR2)
where $e_{k}(\bar{\bm{X}}_{k+1},\bm{W}):=g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})-g_{k}(\bar{\bm{X}}_{k},\bm{W})$ for $k=1,\ldots,K$.
For $k=1,\ldots,K$ define:
$\gamma_{k}(\bar{\bm{X}}_{k},\bm{W}):=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})\mid\bar{\bm{X}}_{k},\bm{W}]=:\mathrm{E}[g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})\mid\bar{\bm{X}}_{k},T=c_{k+1},\bm{W}],$
$\alpha_{k}(\bar{\bm{X}}_{k},\bm{W}):=\prod_{j=1}^{k}\frac{\mathrm{d}P^{(c_{j})}_{X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}}}{\mathrm{d}P^{(c_{k+1})}_{X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}}}(X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}).$
Then (MR2) is equal to the conditional expectation of $Y$ given $\bm{W}$ under the counterfactual distribution $P_{Y\mid\bm{W}}^{(c)}$ if, for every $k=1,\ldots,K$, either $g_{k}(\bar{\bm{X}}_{k},\bm{W})=\gamma_{k}(\bar{\bm{X}}_{k},\bm{W})$ or $a_{k}(\bar{\bm{X}}_{k},\bm{W})=\alpha_{k}(\bar{\bm{X}}_{k},\bm{W})$, but not necessarily both.
The proof is almost identical to that of Theorem 2.5, and therefore we omit it here.
An unconditional version of the above can then be obtained by averaging Equation MR2 over the distribution of the pre-treatment covariates $\bm{W}$. This identification strategy, however, introduces new nuisance functions, the expectations conditional on $\bm{W}$. Unless $\bm{W}$ is discrete with a finite number of values, it will generally not be possible to estimate these at the parametric rate. Hence, below we give an estimating equation that is robust also to the adjustment for covariates. This introduces an additional debiasing term in the equation, the one corresponding to $k=0$ in the sum.
Proposition C.2 (Identification with Covariates, Unconditional Version).
Let $g_{k}(\bar{\bm{X}}_{k},\bm{W})$, $a_{k}(\bar{\bm{X}}_{k},\bm{W})$ be any candidate functions such that $\mathrm{E}[g_{k}(\bar{\bm{X}}_{k},\bm{W})^{2}]<\infty$, $\mathrm{E}[a_{k}(\bar{\bm{X}}_{k},\bm{W})^{2}]<\infty$ for $k=0,\ldots,K$. Consider the following estimating equation:
$\mathrm{E}[g_{0}(\bm{W})]+\sum_{k=0}^{K}\mathrm{E}[a_{k}(\bar{\bm{X}}_{k},\bm{W})e_{k}(\bar{\bm{X}}_{k+1},\bm{W})],$ (MR3)
where $e_{k}(\bar{\bm{X}}_{k+1},\bm{W}):=g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})-g_{k}(\bar{\bm{X}}_{k},\bm{W})$ for $k=1,\ldots,K$.
For $k=1,\ldots,K$ define:
$\gamma_{k}(\bar{\bm{X}}_{k},\bm{W}):=\mathrm{E}_{(c_{k+1})}[g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})\mid\bar{\bm{X}}_{k},\bm{W}]=:\mathrm{E}[g_{k+1}(\bar{\bm{X}}_{k+1},\bm{W})\mid\bar{\bm{X}}_{k},T=c_{k+1},\bm{W}],$
$\alpha_{k}(\bar{\bm{X}}_{k},\bm{W}):=\frac{\bm{1}\{T=c_{k+1}\}}{\Pr(T=c_{k+1}\mid\bm{W})}\prod_{j=1}^{k}\frac{\mathrm{d}P^{(c_{j})}_{X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}}}{\mathrm{d}P^{(c_{k+1})}_{X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}}}(X_{j}\mid\bar{\bm{X}}_{j-1},\bm{W}).$
Moreover, let:
$\gamma_{0}(\bm{W})=\mathrm{E}[g_{1}(X_{1},\bm{W})\mid T=c_{1},\bm{W}],\qquad\alpha_{0}(\bm{W})=\frac{\bm{1}\{T=c_{1}\}}{\Pr(T=c_{1}\mid\bm{W})}.$
Then (MR3) is equal to the (unconditional) expectation of $Y$ under the counterfactual distribution $\int P_{Y\mid\bm{W}}^{(c)}\mathrm{d}P_{\bm{W}}$ if, for every $k=0,\ldots,K$, either $g_{k}(\bar{\bm{X}}_{k},\bm{W})=\gamma_{k}(\bar{\bm{X}}_{k},\bm{W})$ or $a_{k}(\bar{\bm{X}}_{k},\bm{W})=\alpha_{k}(\bar{\bm{X}}_{k},\bm{W})$, but not necessarily both.
Notice that, to adjust for covariates, we need to learn two additional nuisance functions: (i) a final regression step that conditions only on covariates, and (ii) the propensity score, $\Pr(T=1\mid\bm{W})$. These are the same nuisance functions that appear in other problems in causal inference, such as estimating the ATE under unconfoundedness.
Appendix D Simplifications to the Identifying Equation
Remark D.1 (Simplification I: Tower Law).
Recall the Tower Law for conditional expectations:
$\mathrm{E}[\mathrm{E}[A\mid B,C]\mid B]=\mathrm{E}[A\mid B].$
$=\mathrm{E}_{(c_{k})}[\gamma_{k+1}(\bar{\bm{X}}_{k+1})\mid\bar{\bm{X}}_{k-1}].$
Remark D.2 (Simplification II: Conditional Independence).
Suppose that $A\perp B\mid C$:
[…]
Example D.3.
Consider $K=3$ and the causal Markov factorization depicted in Figure 6. (We omit $T$ from the DAG, with the understanding that the distribution of all the variables depends on $T$.) This factorization implies that $X_{2}\perp X_{3}\mid X_{1}$. Suppose we are interested in estimating $\theta^{\bm{c}}$ for $\bm{c}=\langle 1,0,1,0\rangle$. D.2 implies that we can exchange the order of integration of $X_{2}$ and $X_{3}$. Because we’re integrating $X_{1}$ and $X_{3}$ with respect to the same distribution $P_{(\bm{X},Y)}^{(1)}$, and $X_{2}$ and $Y$ with respect to the same distribution $P_{(\bm{X},Y)}^{(0)}$, D.1 helps us in reducing the number of unknown functions to estimate. In particular, we can use the doubly-robust estimating equation:
where the true value of the unknown functions is:
Thus, we only need to estimate 2 unknown functions (rather than 6). ∎
Example D.4.
An important example is when all explanatory variables are independent, i.e., the causal Markov factorization is $P_{(\bm{X},Y)}=P_{Y\mid\bm{X}}\prod_{k=1}^{K}P_{X_{k}}$. For a given change vector $\bm{c}$, let $\bm{X}_{\bm{c}}=(X_{k}:c_{k}=1)$ and $\bm{X}_{-\bm{c}}=(X_{k}:c_{k}=0)$. Thanks to Remarks D.1 and D.2 we can give a doubly-robust estimating equation for $\theta^{\bm{c}}$ that only depends only on 2 unknown functions, regardless of what $K$ is. Whenever $c_{K+1}=0$ (i.e., we want to keep the conditional distribution of $Y\mid\bm{X}$ as in sample 0), the doubly-robust estimating equation will be:
where the true value of the unknown functions is:
where the true value of the unknown functions is:
Appendix E Multiplier Bootstrap
Appendix F Monte-Carlo Simulations
Design 1